# THE POLARIZATION AND MAGNETIC FIELD OF THE RADIO ARC AS OBSERVED BY ALMA AT 100 GHZ

### v1

@@ I INTRODUCTION
The Galactic Center (GC) hosts an extreme environment, with higher densities (103-106 cm-3, Mills2018), temperatures (50-300 K, Krieger2017), magnetic field strengths (100s G-10s mG, e.g., Yusef-Zadeh1987; Chuss2003a; Pillai2015; Guerra2023), and more turbulent velocities ( km s-1 Kauffmann2017) than in the Galactic disk. The GC is the closest galactic nuclear region to Earth at 8.2 kpc (Abuter2019), making it the best available laboratory to study such an environment. Assuming that the GC is representative of other galactic centers, we can infer the properties of extragalactic nuclear regions that are too distant to be resolved with current instruments.
@@
Radio observations towards the GC have revealed a population of thin, glowing strands with lengths up to 10s of parsecs and thicknesses of only a fraction of a parsec (e.g., YMC1984; Yusef-Zadeh2022; Yusef-Zadeh1986a; Morris1996a; Gray1995; Pare2019; Pare2021; Pare2022; Pare2024b). Due to the non-thermal nature of these structures they have become known as the non-thermal filaments (NTFs). These NTFs are highly polarized sources (percentage polarizations 50%) illuminated by synchrotron-emitting relativistic electrons. The magnetic field internal to these structures is generally oriented parallel to the NTFs (YWP1997; Lang1999a; Lang1999b). Since the NTFs are generally oriented perpendicular to the Galactic plane, this suggests a vertical magnetic field that pervades the GC (YMC1984; Yusef-Zadeh1987; LaRosa2006; Morris2006sum; Morris2015). The Radio Arc is one of the brightest systems of NTFs, and was also the first discovered NTF and therefore has been studied across multiple wavelengths at increasingly higher resolution (e.g., YMC1984; Inoue1989; Pare2019).
@@
More recent observations of NTFs have begun to reveal examples where the NTF magnetic field orientation is apparently rotated or even perpendicular to the orientation of the filaments. Recent Very Large Array (VLA) observations at 10 GHz of the Radio Arc and other NTFs have shown alternating magnetic fields that regularly transition from parallel to rotated with respect to the filament orientation (Pare2019; Pare2024b). These rotated field regions may be an intrinsic feature of the Radio Arc or may be caused by interactions with local and foreground magnetized structures (Simpson2007; Butterfield2018). It is also possible that intervening Faraday rotation effects have not been fully corrected (Pare2021). The VLA observations were not able to distinguish between these possible scenarios because of the significance of Faraday effects at 10 GHz.
@@
The 100 GHz frequency of the Atacama Large Millimeter/submillimeter Array (ALMA) observations reported here are much less impacted by Faraday effects because of the dependence of the Faraday effect, allowing us to refine our understanding of whether the orientation offset of the inferred magnetic field with respect to the NTF orientation is a spurious Faraday rotation effect. The Radio Arc is known to be bright at 100 GHz frequencies, as shown from previous 150 GHz single-dish observations made using the Nobeyama 45-m telescope (Reich2000). The Radio Arc is also visible in 2 mm 30-m IRAM observations (Staguhn2019), although neither of these observations have polarimetry. Previous polarimetric observations at 90 GHz also exist for the Radio Arc as obtained using ACTpol (Guan2021), but with a resolution of 1′. The polarimetric aspect of the ALMA observations presented here therefore represent the first time that the magnetic field for this source has been studied at such a high frequency at sub-arcsecond resolution.
@@ II.1 Observations
We observed the Radio Arc in 150 pointings using ALMA Band 3 with a central wavelength of 104.5 GHz (project code: 2024.1.00515.S). The resulting mosaic has a resolution of 0.3″ and the total observation time on-source was 3 hours (three 1-hour sessions), resulting in a sensitivity level of 100 Jy. These ALMA observations have baseline lengths ranging from 15 to 260 meters. More details regarding the observations are displayed in Table 1.
@@ II.2 Calibration
The data reduction was done using the Common Astronomy Software Application (CASA) utilizing the ALMA calibration pipeline (CASA pipeline version 6.6.6.17), which now includes polarization calibration. The pipeline calibrated products were of sufficiently high quality to not require any additional manual calibration. We verified that J1924-2914 was observed to be significantly polarized by imaging the polarization calibrator after applying the pipeline calibration to the data set.
@@ II.3 Imaging
The 100 GHz observations presented here represent the first time ALMA has been used to observe the unique NTF structures in the GC polarimetrically at such high resolution. Figure 1 displays the total intensity and polarized intensity of the Radio Arc obtained from ALMA using the calibration pipeline. The top panel of Figure 1 marks the locations of the unusual N3 point source located within the NTFs (Ludovici2016) and the nearby Pistol H II region (YMC1984; Figer1998; Figer1999pist).
@@
We also made spectro-polarimetric cubes with 4 channels, one channel for each of the four spectral windows comprising the ALMA observations. This spectral cube allowed for an assessment of the significance of potential Faraday effects by measuring the Rotation Measure (RM) towards the Radio Arc such that: , where is the observed polarization angle (rad), is the Rotation Measure observed towards the Radio Arc (rad m-2), is the wavelength of the observation, and is the intrinsic polarization angle emitted by the source. By fitting the spectrum of the polarization angle as a function of wavelength squared with a linear model for each line of sight we obtain the the RM as the slope of the best-fit line for each pixel. We found from this model fitting that RM effects are negligible at the 100 GHz ALMA frequencies of this data set.
@@ III.1 Debiased Polarized Intensity and the Derived Magnetic Field
We also observe significant polarized intensity only in the Southernmost extent of the Radio Arc filaments. The Northern portion of the Radio Arc exhibits no significant polarization. This depolarization effect has been observed in previous polarimetric observations of the Radio Arc, with the depolarization possibly being the result of a more complicated line-of-sight geometry in this region of the Radio Arc or the increased proximity with prominent H II regions (Yusef-Zadeh1987; Pare2019).
@@ IV.1 Polarized Intensity Features
As with previously observed NTFs, our ALMA observations reveal a discontinuous or “patchy” polarized intensity distribution that generally traces the NTF structure (e.g., Gray1995; Lang1999a; Pare2019; Pare2024b). We only observe significant polarized intensity in Southern declinations of the Radio Arc as can be seen in Figures 1 and 2. The depolarization of the Radio Arc in its Northern extent has been observed in lower-frequency (10 GHz) observations previously (e.g., Yusef-Zadeh1987; Inoue1989). This depolarization trend has been attributed to beam depolarization, line-of-sight contamination from molecular clouds coinciding with the Northern extent of the Radio Arc causing cancellation of the polarized signal of the Radio Arc, or the prominent H II regions in the Northern region of the Radio Arc (Yusef-Zadeh1988). Since this depolarization is still observed at the 100 GHz ALMA observations presented here with higher resolution, we favor the possibilities that either the H II regions or the complex molecular structures in the Northern portion of the Radio Arc could be producing the depolarization observed in this region of the NTF.
@@
To assess the origin of the polarization we note that previous VLA observations of the Radio Arc have revealed polarized intensity features that lack a significant total intensity counterpart (Inoue1989; Pare2019). These polarized intensity extensions are all observed to have fractional polarization p = P/I 1.0, indicative of missing total intensity as a result of the interferometric properties of the observations. We find an average fractional polarization of 1.4 for PF2. This result is non-physical, as one would not expect to obtain a fractional polarization 1.0; however, this result supports the possibility that there is a total intensity counterpart to this polarization structure that is not sampled by the interferometric ALMA observations. This conclusion is corroborated by the fact that PF1, which coincides with a total intensity NTF counterpart, has an average fractional polarization of 0.7.
@@
There is evidence supporting both of the above scenarios: previous VLA observations indicate that a radio shell of emission could be impacting the polarimetric distributions obtained for the Radio Arc (Pare2019; Pare2021). This brightness of the shell only appears on the limb of the shell projected in the plane of the sky because it is a larger structure than the maximum recoverable scales of the VLA observations. This same effect, however, does reveal the density enhancement within the shell. Observational evidence for the Radio Shell is also readily apparent in other interferometric observations of the region made using the VLA and MeerKAT (Lang2010; Heywood2022).
The radio shell has also been observed at infrared wavelengths (Egan1998; Levine1999; Price2001; Simpson2007), and these infrared observations reveal the structure to have a diameter of 10′. This shell encompasses the Quintuplet star cluster, which led Figer1999 to postulate that this shell could be a shock front generated by winds from this star cluster. Pare2021 argue that their Faraday effect analysis, coupled with these infrared observations, supports the possibility that the density enhancement of the shell alternately lies in the foreground and background along the line-of-sight with respect to the Radio Arc filaments.
However, the fact that the polarization of PF2 follows the NTF filamentary direction (as seen in Figure 2) supports the possibility of the centrally concentrated polarization. Furthermore, the ALMA observations analyzed here have a higher resolution (0.3″) compared to the previous VLA observations (1.0″), indicating that the ALMA observations could be resolving more of the fine-scale features of the Radio Arc polarization.
@@ IV.2 Radio Arc Magnetic Field from ALMA
The observed magnetic field inferred from ALMA has two enhancements in orientation as seen in the lower panel of Figure 3, with the enhancement for PF1 corresponding to angles ranging from 20 to 60∘ whereas the enhancement for PF2 corresponds to angles ranging from -80 to -20∘. The systematically different orientations between PF1 and PF2, coupled with the uniform field orientation within each filament, lies in stark contrast to what has been observed previously at lower frequencies (10 GHz) where the magnetic field is predominantly parallel to the Radio Arc filament orientation (Yusef-Zadeh1987a; Inoue1989; Pare2019, e.g.,). There are regions within the Radio Arc where a rotated magnetic field orientation was observed with 10 GHz VLA data (Pare2019). The locations of these rotated magnetic field regions as observed by the VLA are marked with blue rectangles in the top panel of Figure 3. We find that the magnetic field orientations of PF1 and PF2 do not generally align with the rotated regions observed in the VLA.
The alternating magnetic field pattern observed from the VLA was argued to originate from electron density variations in a shell of radio emission that appears to envelop the Radio Arc (Pare2019; Pare2021), where the rotated magnetic field regions coincide with higher electron density regions of the radio shell that lie foreground to the Radio Arc along the line-of-sight. This geometry is depicted as a sketch in the left-hand panel of Figure 4 (panel a) where the Radio Arc filaments are shown as black lines and the radio shell is shown as the colored circle. The density enhancements are depicted as the dashed rectangles within the circle where one of the enhancements is foreground to the Radio Arc along the line-of-sight, resulting in a rotated magnetic field orientation. The other density enhancement is background to the NTFs, tracing the portion of the shell that is on the opposite side of the Radio Arc from our direction of observation. The background density enhancement therefore does not induce a rotated magnetic field geometry. This sketch was motivated by comparisons to the polarized intensity, rotation measure, and magnetic field distributions studied by Pare2019 and was subsequently corroborated using a more in-depth analysis of the line-of-sight Faraday effects encountered towards the source in Pare2021.
@@
We note, however, that the previous VLA observations also identified total intensity features that could possibly be density enhancements of the Radio Shell. These features are labeled in Figure 1 of Pare2019 as “helical segments,” and the coincidence of these features with the rotated magnetic field regions is used to support the sketch shown in the left-hand panel of Figure 4. Indeed, if the large fractional polarizations of PF2 are a result of centrally concentrated polarization within the Radio Arc filaments, as discussed in Section IV.1, then the uniform (but rotated) magnetic fields observed from these ALMA observations could indicate variations in the magnetic field between different Radio Arc filaments, rather than effects caused by a foreground intervening medium. This scenario is depicted in the right-hand panel of Figure 4 (panel c), where the central polarized cores are represented as green lines and the enveloping Radio Shell has the density enhancements inferred from the previous VLA observations (Pare2019; Pare2021).
We argue that panels b and c of Figure 4 are more accurate representations of the geometry local to the Radio Arc. Based on the ALMA observations presented here and the previous VLA observations, we slightly prefer the sketch shown in panel c of Figure 4 as the most accurate representation of the geometry local to the Radio Arc: the higher resolution ALMA observations are not as sensitive to the effects caused by the density enhancements of the Radio Shell, but are more sensitive to the potential of the polarization being centrally concentrated within the filaments. Conversely, the previous 10 GHz VLA observations are preferentially effected in the opposite manner, resulting in the alternating magnetic field pattern inferred from those observations (Pare2019; Pare2021).
Finally, we study whether there are any deformations in the Radio Arc that could help determine possible interaction between the NTFs and the Radio Shell. Tsuboi1997 use deformations in the Southernmost extent of the Radio Arc to infer interaction between the Radio Arc and a shell traced by CS 1 – 0 and CS 2 – 1 emission (this is a distinct shell from the one discussed in this present work). This Southern extent of the Radio Arc is outside the ALMA field of view, but we can inspect the observed portion of the Radio Arc for deformations to infer possible interaction. We do not see any perturbations, but we note that this could be a result of the Radio Shell not having enough momentum to induce such perturbations given the high rigidity of the NTFs. Furthermore, the likely high magnetic field strengths of the Radio Arc NTFs (with estimates being 100s of G) would mean that such perturbations would propagate quickly through the NTFs. The lack of deformations here can therefore not be invoked to conclusively rule out the possibility that these structures are interacting.

### v2

@@ I INTRODUCTION
The Galactic Center (GC) hosts an extreme environment, with higher densities (103-106 cm-3, Mills et al., 2018), temperatures (50-300 K, Krieger et al., 2017), magnetic field strengths (100s G-10s mG, e.g., Yusef-Zadeh and Morris, 1987a; Chuss et al., 2003; Pillai et al., 2015; Guerra et al., 2023), and more turbulent velocities ( km s-1 Kauffmann et al., 2017) than in the Galactic disk. The GC is the closest galactic nuclear region to Earth at 8.2 kpc (Gravity Collaboration et al., 2019), making it the best available laboratory to study such an environment. Assuming that the GC is representative of other galactic centers, we can infer the properties of extragalactic nuclear regions that are too distant to be resolved with current instruments.
@@
Radio observations towards the GC have revealed a population of thin, glowing strands with lengths up to 10s of parsecs and thicknesses of only a fraction of a parsec (e.g., Yusef-Zadeh et al., 1984, 2022, 1986; Morris, 1996; Gray et al., 1995; Paré et al., 2019, 2021, 2022, 2024). Due to the non-thermal nature of these structures they have become known as the non-thermal filaments (NTFs). These NTFs are highly polarized sources (percentage polarizations 50%) illuminated by synchrotron-emitting relativistic electrons. The magnetic field internal to these structures is generally oriented parallel to the NTFs (Yusef-Zadeh et al., 1997; Lang et al., 1999a, b). Since the NTFs are generally oriented perpendicular to the Galactic plane, this suggests a vertical magnetic field that pervades the GC (Yusef-Zadeh et al., 1984; Yusef-Zadeh and Morris, 1987a; La Rosa et al., 2006; Morris, 2006, 2015). The Radio Arc is one of the brightest systems of NTFs, and was also the first discovered NTF and therefore has been studied across multiple wavelengths at increasingly higher resolution (e.g., Yusef-Zadeh et al., 1984; Inoue et al., 1989; Paré et al., 2019).
@@
More recent observations of NTFs have begun to reveal examples where the NTF magnetic field orientation is apparently rotated or even perpendicular to the orientation of the filaments. Recent Very Large Array (VLA) observations at 10 GHz of the Radio Arc and other NTFs have shown alternating magnetic fields that regularly transition from parallel to rotated with respect to the filament orientation (Paré et al., 2019, 2024). These rotated field regions may be an intrinsic feature of the Radio Arc or may be caused by interactions with local and foreground magnetized structures (Simpson et al., 2007; Butterfield et al., 2018). It is also possible that intervening Faraday rotation effects have not been fully corrected (Paré et al., 2021). The VLA observations were not able to distinguish between these possible scenarios because of the significance of Faraday effects at 10 GHz.
@@
The 100 GHz frequency of the Atacama Large Millimeter/submillimeter Array (ALMA) observations reported here are much less impacted by Faraday effects because of the dependence of the Faraday effect, allowing us to refine our understanding of whether the orientation offset of the inferred magnetic field with respect to the NTF orientation is a spurious Faraday rotation effect. The Radio Arc is known to be bright at 100 GHz frequencies, as shown from previous 150 GHz single-dish observations made using the Nobeyama 45-m telescope (Reich et al., 2000). The Radio Arc is also visible in 2 mm 30-m IRAM observations (Staguhn et al., 2019), although neither of these observations have polarimetry. Previous polarimetric observations at 90 GHz also exist for the Radio Arc as obtained using ACTpol (Guan et al., 2021), but with a resolution of 1′. The polarimetric aspect of the ALMA observations presented here therefore represent the first time that the magnetic field for this source has been studied at such a high frequency at sub-arcsecond resolution.
@@ II.1 Observations
We observed the Radio Arc in 150 pointings using ALMA Band 3 with a central wavelength of 104.5 GHz (project code: 2024.1.00515.S). The resulting mosaic has a resolution of 1.5″ with a pixel size of 0.3″ (for 5 pixels across each beam) and the total observation time on-source was 3 hours (three 1-hour sessions), resulting in a sensitivity level of 100 Jy. These ALMA observations have baseline lengths ranging from 15 to 260 meters. More details regarding the observations are displayed in Table 1.
@@ II.2 Calibration
The data reduction was done using the Common Astronomy Software Application (CASA) utilizing the ALMA calibration pipeline (CASA pipeline version 6.6.6.17), which now includes polarization calibration. The pipeline calibrated products were of sufficiently high quality to not require any additional manual calibration. We verified that J1924-2914 was observed to be significantly polarized by imaging the polarization calibrator after applying the pipeline calibration to the data set. We obtained a Stokes intensity of 5 Jy beam-1 and Stokes and intensities of 0.1 Jy beam-1, equating to a percentage polarization of 2% for J1924-2914. This Stokes value and percentage polarization agree with previous ALMA observations of this calibrator (e.g., Goddi et al., 2021).
@@ II.3 Imaging
The 100 GHz observations presented here represent the first time ALMA has been used to observe the unique NTF structures in the GC polarimetrically at such high resolution. Figure 1 displays the total intensity and polarized intensity of the Radio Arc obtained from ALMA using the calibration pipeline. The top panel of Figure 1 marks the locations of the unusual N3 point source located within the NTFs (Ludovici et al., 2016) and the nearby Pistol H II region (Yusef-Zadeh et al., 1984; Figer et al., 1998, 1999b).
@@
We also made spectro-polarimetric cubes with 4 channels, one channel for each of the four spectral windows comprising the ALMA observations. This spectral cube allowed for an assessment of the significance of potential Faraday effects by measuring the Rotation Measure (RM) towards the Radio Arc such that: , where is the observed polarization angle (rad), is the Rotation Measure observed towards the Radio Arc (rad m-2), is the wavelength of the observation, and is the intrinsic polarization angle emitted by the source. By fitting the spectrum of the polarization angle as a function of wavelength squared with a linear model for each line of sight we obtain the the RM as the slope of the best linear fit for each pixel. We found from this model fitting that RM effects are negligible (10 rad m 5 rad m-2) at the 100 GHz ALMA frequencies of this data set for the lines of sight corresponding to the significant polarization studied. Rotation Measures of this magnitude will not significantly rotate the observed polarization angle at the wavelengths of this ALMA observation.
@@ III.1 Debiased Polarized Intensity and the Derived Magnetic Field
We also observe significant polarized intensity only in the Southernmost extent of the Radio Arc filaments. The Northern portion of the Radio Arc exhibits no significant polarization. This depolarization effect has been observed in previous polarimetric observations of the Radio Arc, with the depolarization possibly being the result of a more complicated line-of-sight geometry in this region of the Radio Arc or the increased proximity with prominent H II regions (Yusef-Zadeh and Morris, 1987a; Paré et al., 2019).
@@ IV.1 Polarized Intensity Features
As with previously observed NTFs, our ALMA observations reveal a discontinuous or “patchy” polarized intensity distribution that generally traces the NTF structure (e.g., Gray et al., 1995; Lang et al., 1999a; Paré et al., 2019, 2024). We only observe significant polarized intensity in Southern declinations of the Radio Arc as can be seen in Figures 1 and 2. The depolarization of the Radio Arc in its Northern extent has been observed in lower-frequency (10 GHz) observations previously (e.g., Yusef-Zadeh and Morris, 1987a; Inoue et al., 1989). This depolarization trend has been attributed to beam depolarization, line-of-sight contamination from molecular clouds coinciding with the Northern extent of the Radio Arc causing cancellation of the polarized signal of the Radio Arc, or the prominent H II regions in the Northern region of the Radio Arc (Yusef-Zadeh and Morris, 1988). Since this depolarization is still observed at the 100 GHz ALMA observations presented here with higher resolution, we favor the possibilities that either the H II regions or the complex molecular structures in the Northern portion of the Radio Arc could be producing the depolarization observed in this region of the NTF.
@@
To assess the origin of the polarization we note that previous VLA observations of the Radio Arc have revealed polarized intensity features that lack a significant total intensity counterpart (Inoue et al., 1989; Paré et al., 2019). These polarized intensity extensions are all observed to have fractional polarization p = P/I 1.0, indicative of missing total intensity as a result of the interferometric properties of the observations. We find an average fractional polarization of 1.4 for PF2. This result is non-physical, as one would not expect to obtain a fractional polarization 1.0; however, this result supports the possibility that there is a total intensity counterpart to this polarization structure that is not sampled by the interferometric ALMA observations. The C43-3 ALMA array configuration has a maximum recoverable scale of 16.2″. Coherent structures much larger than this limit, such as the total intensity counterpart of PF2, are not being sampled. This conclusion is corroborated by the fact that PF1, which coincides with a total intensity NTF counterpart, has an average fractional polarization of 0.7.
@@
There is evidence supporting both of the above scenarios: previous VLA observations indicate that a radio shell of emission could be impacting the polarimetric distributions obtained for the Radio Arc (Paré et al., 2019, 2021). This brightness of the shell only appears on the limb of the shell projected in the plane of the sky because at an angular size of 10′it is a larger structure than the maximum recoverable scales of the VLA observations (Simpson et al., 2007). This same edge-brightening effect, however, does reveal the density enhancement within the shell. Observational evidence for the Radio Shell is also readily apparent in other interferometric observations of the region made using the VLA and MeerKAT (Lang et al., 2010; Heywood et al., 2022).
The radio shell has also been observed at infrared wavelengths (Egan et al., 1998; Levine et al., 1999; Price et al., 2001; Simpson et al., 2007), and these infrared observations reveal the structure to have a diameter of 10′. This shell encompasses the Quintuplet star cluster, which led Figer et al. (1999a) to postulate that this shell could be a shock front generated by winds from this star cluster. Paré et al. (2021) argue that their Faraday effect analysis, coupled with these infrared observations, supports the possibility that the density enhancement of the shell alternately lies in the foreground and background along the line-of-sight with respect to the Radio Arc filaments.
However, the fact that the polarization of PF2 follows the NTF filamentary direction (as seen in Figure 2) supports the possibility of the centrally concentrated polarization. Furthermore, the ALMA observations analyzed here have a higher resolution (pixel size of 0.3″) compared to the previous VLA observations (1.0″), indicating that the ALMA observations could be resolving more of the fine-scale features of the Radio Arc polarization.
@@ IV.2 Radio Arc Magnetic Field from ALMA
The observed magnetic field inferred from ALMA has two enhancements in orientation as seen in the lower panel of Figure 3, with the enhancement for PF1 corresponding to angles ranging from 20 to 60∘ whereas the enhancement for PF2 corresponds to angles ranging from -80 to -20∘. We report an uncertainty on the magnetic field angles of which was determined by propagating the uncertainties obtained on the Stokes and distributions. The systematically different orientations between PF1 and PF2, coupled with the uniform field orientation within each filament, lies in stark contrast to what has been observed previously at lower frequencies (10 GHz) where the magnetic field is predominantly parallel to the Radio Arc filament orientation (Yusef-Zadeh and Morris, 1987b; Inoue et al., 1989; Paré et al., 2019, e.g.,). There are regions within the Radio Arc where a rotated magnetic field orientation was observed with 10 GHz VLA data (Paré et al., 2019). The locations of these rotated magnetic field regions as observed by the VLA are marked with blue rectangles in the top panel of Figure 3. We find that the magnetic field orientations of PF1 and PF2 do not generally align with the rotated regions observed in the VLA.
The alternating magnetic field pattern observed from the VLA was argued to originate from electron density variations in a shell of radio emission that appears to envelop the Radio Arc (Paré et al., 2019, 2021), where the rotated magnetic field regions coincide with higher electron density regions of the radio shell that lie foreground to the Radio Arc along the line-of-sight. This geometry is depicted as a sketch in the left-hand panel of Figure 4 (panel a) where the Radio Arc filaments are shown as black lines and the radio shell is shown as the colored circle. The density enhancements are depicted as the dashed rectangles within the circle where one of the enhancements is foreground to the Radio Arc along the line-of-sight, resulting in a rotated magnetic field orientation. The other density enhancement is background to the NTFs, tracing the portion of the shell that is on the opposite side of the Radio Arc from our direction of observation. The background density enhancement therefore does not induce a rotated magnetic field geometry. This sketch was motivated by comparisons to the polarized intensity, rotation measure, and magnetic field distributions studied by Paré et al. (2019) and was subsequently corroborated using a more in-depth analysis of the line-of-sight Faraday effects encountered towards the source in Paré et al. (2021).
@@
We note, however, that the previous VLA observations also identified total intensity features that could possibly be density enhancements of the Radio Shell. These features are labeled in Figure 1 of Paré et al. (2019) as “helical segments,” and the coincidence of these features with the rotated magnetic field regions is used to support the sketch shown in the left-hand panel of Figure 4. Indeed, if the large fractional polarizations of PF2 are a result of centrally concentrated polarization within the Radio Arc filaments, as discussed in Section IV.1, then the uniform (but rotated) magnetic fields observed from these ALMA observations could indicate variations in the magnetic field between different Radio Arc filaments, rather than effects caused by a foreground intervening medium. This scenario is depicted in the right-hand panel of Figure 4 (panel c), where the central polarized cores are represented as green lines and the enveloping Radio Shell has the density enhancements inferred from the previous VLA observations (Paré et al., 2019, 2021).
We argue that panels b and c of Figure 4 are more accurate representations of the geometry local to the Radio Arc. Based on the ALMA observations presented here and the previous VLA observations, we slightly prefer the sketch shown in panel c of Figure 4 as the most accurate representation of the geometry local to the Radio Arc: the higher resolution ALMA observations are not as sensitive to the effects caused by the density enhancements of the Radio Shell, but are more sensitive to the potential of the polarization being centrally concentrated within the filaments. Conversely, the previous 10 GHz VLA observations are preferentially effected in the opposite manner, resulting in the alternating magnetic field pattern inferred from those observations (Paré et al., 2019, 2021).
Finally, we study whether there are any deformations in the Radio Arc that could help determine possible interaction between the NTFs and the Radio Shell. Tsuboi et al. (1997) use deformations in the Southernmost extent of the Radio Arc to infer interaction between the Radio Arc and a shell traced by CS 1 – 0 and CS 2 – 1 emission (this is a distinct shell from the one discussed in this present work). This Southern extent of the Radio Arc is outside the ALMA field of view, but we can inspect the observed portion of the Radio Arc for deformations to infer possible interaction. We do not see any such perturbations, but we note that this could be a result of the Radio Shell not having enough momentum to induce such perturbations given the high rigidity of the NTFs. Furthermore, the likely high magnetic field strengths of the Radio Arc NTFs (with estimates being 100s of G) would mean that such perturbations would propagate quickly through the NTFs as Alfvén waves. The lack of deformations here can therefore not be invoked to conclusively rule out the possibility that these structures are interacting.